# Multi-resolution filters via linear projection for large spatio-temporal datasets

### v2

@@ 1 Introduction
This paper proposes a multi-resolution filter via linear projection (MRF-lp) for large spatio-temporal datasets. This proposed method approximates the forecast covariance matrix in the Kalman filter by using a multi-resolution approximation via linear projection (MRA-lp) (Hirano,, 2021) which captures both the large- and small-scale spatial variations, unlike the covariance tapering (e.g., Furrer et al.,, 2006; Hirano and Yajima,, 2013) and low rank approaches (e.g., Banerjee et al.,, 2008). In addition, as a remarkable property, the MRF-lp preserves a block sparse structure of some matrices through time. Consequently, we can quickly calculate a good approximation of the filtering mean for large spatio-temporal datasets. Our proposed MRF-lp can be regarded as an extension of the MRA-lp to spatio-temporal datasets and a generalization of a multi-resolution filter (MRF) developed by Jurek and Katzfuss, (2021). As a result, the aforementioned desirable properties of the MRF-lp are inherited from the MRA-lp and MRF. Moreover, building upon the idea of Zilber and Katzfuss, (2021) and Jurek and Katzfuss, (2022), we also discuss extensions of the MRF-lp to a nonlinear and non-Gaussian case where the system model is nonlinear, and the observation model is a distribution from an exponential family. The extensions are achieved by using the extended Kalman filter via the first-order Taylor expansion, the Laplace approximation, and the Newton–Raphson method.
The rest of this paper is organized as follows. We review linear Gaussian state-space models and the Kalman filter for spatio-temporal datasets in Section 2. In Section 3, we introduce the MRA-lp for spatio-temporal state-space models. Section 4 describes our proposed MRF-lp. In Section 5, we discuss the extension to nonlinear and non-Gaussian state-space models. In Section 6, we illustrate the results of the simulation studies and real data analysis. Finally, we conclude this study in Section 7. The appendices contain the derivations of some algorithms, the technical proofs of the propositions, and the EnKF algorithm in the nonlinear and Gaussian case. In the supplementary material, we provide the overview of the MRA-lp, the derivations of the linear and nonlinear transformations in the system model and the detailed conditions in the simulation studies and real data analysis, and the additional simulation study. To describe a physical evolution precisely and attain a sophisticated forecast, these transformations are derived from two kinds of partial differential equations in natural science.
@@ 3 Multi-resolution approximation via linear projection for spatio-temporal state-space models
Finally, we define an $r^{\prime}_{j_{1},\ldots,j_{m}}\times r_{j_{1},\ldots,j_{m}}$ matrix $\Phi_{j_{1},\ldots,j_{m}}$ ($m=0,\ldots,M$, $1\leq r^{\prime}_{j_{1},\ldots,j_{m}}\leq r_{j_{1},\ldots,j_{m}}$) where $\operatorname{rank}\left(\Phi_{j_{1},\ldots,j_{m}}\right)=r^{\prime}_{j_{1},\ldots,j_{m}}$, and the row-norm of $\Phi_{j_{1},\ldots,j_{m}}$ is equal to 1. To avoid the computational burden, it is posited that $r^{\prime}_{j_{1},\ldots,j_{m}}\ll n_{\mathcal{G}}$. In Section 3.2, we will discuss how to select $\Phi_{j_{1},\ldots,j_{m}}$ in this study.
Section S1 provides an overview of the MRA-lp where we also explain the role of $\Phi_{j_{1},\ldots,j_{m}}$.
@@ 6.1 Simulation study
Through simulation studies, we evaluated the performances of Algorithms 35. We used Algorithms 35 with $\Phi_{j_{1},\ldots,j_{m}}=\bm{\mathrm{I}}_{r_{j_{1},\ldots,j_{m}}}$ as the MRF, meaning $r_{j_{1},\ldots,j_{m}}=r_{j_{1},\ldots,j_{m}}^{\prime}$ in the MRF. The original methods in the non-Gaussian case correspond to Algorithms 4 and 5 where Algorithm 2 is not applied. Additionally, we conducted a comparison with the EnKF. In Section S4, we evaluated the numerical stability of the MRF-lp in calculating the inverse matrices by using $\Phi_{j_{1},\ldots,j_{m}}$.
@@ 7 Conclusion and future studies
This study proposed the MRF-lp, which can quickly compute the filtering mean for linear Gaussian state-space models. Furthermore, we have extended our proposed method to nonlinear and non-Gaussian state-space models on the basis of Zilber and Katzfuss, (2021) and Jurek and Katzfuss, (2022). The block sparse structure resulting from the MRA-lp makes the MRF-lp the efficient filtering algorithm. The MRF-lp can be regarded as an extension of the MRA-lp to the spatio-temporal datasets and a generalization of the MRF. By increasing $r_{j_{1},\ldots,j_{m}}$ and/or $r^{\prime}_{j_{1},\ldots,j_{m}}$ adequately, the approximation accuracy of the filtering mean can be improved while maintaining the relatively small computational time. Our proposed MRF-lp has the advantage of adjusting $r_{j_{1},\ldots,j_{m}}$ and $r^{\prime}_{j_{1},\ldots,j_{m}}$ separately, unlike the MRF where $r_{j_{1},\ldots,j_{m}}=r^{\prime}_{j_{1},\ldots,j_{m}}$. In the simulation studies and real data analysis, the MRF-lp was generally more efficient than the MRF in terms of approximation accuracy and computational time.
There are some issues to be tackled in the future. First, Jurek and Katzfuss, (2021, 2022) proposed the fast computation algorithm of Rao–Blackwellized particle filter (Doucet and Johansen,, 2011) to estimate parameters in state-space models (e.g., range parameter in the covariance function). Similarly, our proposed MRF-lp would be extended to the particle version. Second, it is difficult for the extended Kalman filter to capture the strongly nonlinear evolution. Based on Chakraborty and Katzfuss, (2022), combining the MRF-lp with the unscented Kalman filter is interesting. Third, similar to
Jurek and Katzfuss, (2023), the fast smoothing algorithm by using the MRA-lp is considered. As shown in Figure 6 of Section 6.2, the filtering distribution degenerates when few observations are obtained at early time points. However, the smoothing distribution may provide a plausible prediction surface. Fourth, the MRF-lp has a lot of tuning parameters such as the highest resolution $M$, the partition number $J_{m}$ of each subregion, the number $r_{j_{1},\ldots,j_{m}}$ of knots, and the rank $r_{j_{1},\ldots,j_{m}}^{\prime}$ of $\Phi_{j_{1},\ldots,j_{m}}$. A practical guideline for choosing them would be helpful. Finally, the general Vecchia approximation proposed by Katzfuss and Guinness, (2021) includes some existing fast computation methods used in spatial statistics and has shown high usefulness in a variety of situations (e.g., Katzfuss et al., 2020a, ; Zilber and Katzfuss,, 2021; Katzfuss et al.,, 2022; Jurek and Katzfuss,, 2022, 2023). It is left for a future study to explore the relationship between the MRF-lp and the fast filtering method in Jurek and Katzfuss, (2022), which adopts a hierarchical Vecchia approximation.

### v3

@@ 1 Introduction
This paper proposes a multi-resolution filter via linear projection (MRF-lp) for large spatio-temporal datasets. This proposed method approximates the forecast covariance matrix in the Kalman filter by using a multi-resolution approximation via linear projection (MRA-lp) (Hirano,, 2021) which captures both the large- and small-scale spatial variations, unlike the covariance tapering (e.g., Furrer et al.,, 2006; Hirano and Yajima,, 2013) and low rank approaches (e.g., Banerjee et al.,, 2008). In addition, as a remarkable property, the MRF-lp preserves a block sparse structure of some matrices through time. Consequently, we can quickly calculate a good approximation of the filtering mean for large spatio-temporal datasets. Our proposed MRF-lp can be regarded as an extension of the MRA-lp to spatio-temporal datasets and a generalization of a multi-resolution filter (MRF) developed by Jurek and Katzfuss, (2021). Additionally, unlike the MRF, the MRF-lp can empirically reduce ill-conditioning of certain matrices and increase numerical stability. The aforementioned desirable properties of the MRF-lp are inherited from the MRA-lp and MRF. Moreover, building upon the idea of Zilber and Katzfuss, (2021) and Jurek and Katzfuss, (2022), we also discuss extensions of the MRF-lp to a nonlinear and non-Gaussian case where the system model is nonlinear, and the observation model is a distribution from an exponential family. The extensions are achieved by using the extended Kalman filter via the first-order Taylor expansion, the Laplace approximation, and the Newton–Raphson method.
The rest of this paper is organized as follows. We review linear Gaussian state-space models and the Kalman filter for spatio-temporal datasets in Section 2. In Section 3, we introduce the MRA-lp for spatio-temporal state-space models. Section 4 describes our proposed MRF-lp. In Section 5, we discuss the extension to nonlinear and non-Gaussian state-space models. In Section 6, we illustrate the results of the simulation studies and real data analysis. Finally, we conclude this study in Section 7. The appendices contain the derivations of some algorithms, the technical proofs of the propositions, and the EnKF algorithm in the nonlinear and Gaussian case. In the supplementary material, we provide the derivations of the linear and nonlinear transformations in the system model and the detailed conditions in the simulation studies and real data analysis. To describe a physical evolution precisely and attain a sophisticated forecast, these transformations are derived from two kinds of partial differential equations in natural science.
@@ 3 Multi-resolution approximation via linear projection for spatio-temporal state-space models
Finally, we define an $r^{\prime}_{j_{1},\ldots,j_{m}}\times r_{j_{1},\ldots,j_{m}}$ matrix $\Phi_{j_{1},\ldots,j_{m}}$ ($m=0,\ldots,M$, $1\leq r^{\prime}_{j_{1},\ldots,j_{m}}\leq r_{j_{1},\ldots,j_{m}}$) where $\operatorname{rank}\left(\Phi_{j_{1},\ldots,j_{m}}\right)=r^{\prime}_{j_{1},\ldots,j_{m}}$, and the row-norm of $\Phi_{j_{1},\ldots,j_{m}}$ is equal to 1. To avoid the computational burden, it is posited that $r^{\prime}_{j_{1},\ldots,j_{m}}\ll n_{\mathcal{G}}$. Intuitively, the role of $\Phi_{j_{1},\ldots,j_{m}}$ is to generate $r^{\prime}_{j_{1},\ldots,j_{m}}$ new compressed knots from $r_{j_{1},\ldots,j_{m}}$ original knots. As shown in the simulation studies of Banerjee et al., (2013) and Hirano, (2021), the numerical stability of the predictive process (Banerjee et al.,, 2008) and multi-resolution approximation (MRA) (Katzfuss,, 2017) can be empirically improved by replacing the original knots with the compressed knots. Likewise, the MRF-lp with the compressed knots is numerically more stable than the MRF with the original knots (see the sixth simulation study in Section 6.1 for details). Other merits of introducing $\Phi_{j_{1},\ldots,j_{m}}$ are described, for example, in Section 3.1. In Section 3.3, we will discuss how to select $\Phi_{j_{1},\ldots,j_{m}}$ in this study.
3.1 Overview of the MRA-lp
In this section, based on Hirano, (2021), we will summarize the MRA-lp. To begin with, we introduce some notations that are needed except for those before Section 3.1. However, note that the new additional notations are available only in Section 3.1 and are sometimes introduced with different definitions after the next subsection. Let $Y_{0}(\bm{s})\sim\mbox{GP}(0,C_{0})$ be a zero-mean Gaussian process with a covariance function $C_{0}(\bm{s},\bm{s}^{*})$ ($\bm{s}$, $\bm{s}^{*}\in D_{0}$ ). Also, a set of knots on each subregion $D_{j_{1},\ldots,j_{m}}$ is denoted by $Q_{j_{1},\ldots,j_{m}}$ ($m=0,\ldots,M$). The knot means the location on $D_{j_{1},\ldots,j_{m}}$. In addition, for a generic Gaussian process $Y(\bm{s})\sim\mbox{GP}(0,C^{\prime})$ and sets of the vectors, that is, $A^{\prime}=\{\bm{a}_{1},\ldots,\bm{a}_{N}\}$ and $B^{\prime}=\{\bm{b}_{1},\ldots,\bm{b}_{M}\}$ ($\bm{a}_{i},\bm{b}_{j}\in\mathbb{R}^{d^{\prime}}$, $i=1,\ldots,N,\;j=1,\ldots,M$), we write $\bm{Y}(A^{\prime})=\left(Y(\bm{a}_{1}),\ldots,Y(\bm{a}_{N})\right)^{\top}$ and $(C^{\prime}(A^{\prime},B^{\prime}))_{ij}=C^{\prime}(\bm{a}_{i},\bm{b}_{j})$ $(i=1,\ldots,N,\;j=1,\ldots,M)$.
$\tau_{0}(\bm{s})$ $=E[Y_{0}(\bm{s})|\Phi_{0}\bm{Y}_{0}(Q_{0})]$
Then, it follows that
where $\widehat{V}_{0}^{0}=\Phi_{0}V_{0}^{0}\Phi_{0}^{\top}$ and $V_{0}^{0}=C_{0}(Q_{0},Q_{0})$. The linear projection uses $C_{\tau_{0}}$ as an approximation of $C_{0}$ for fast computation in large spatial datasets and is identical with the predictive process proposed by Banerjee et al., (2008) in the case of $\Phi_{0}=\bm{\mbox{I}}_{\left|Q_{0}\right|}$.
In the MRA-lp, the calculation of $C_{\tau_{0}}$ is regarded as the linear projection at resolution 0. Based on Katzfuss, (2017), the MRA-lp implements the linear projection on $D_{j_{1},\ldots,j_{m}}$ for the approximation error of $C_{0}$ at each $m$th resolution ($m=1,\ldots,M$). Specifically, for $\bm{s}\in D_{0}$, let $\delta_{m}(\bm{s})\sim\mbox{GP}(0,C_{m})$ ($m=1,\ldots,M$) be a zero-mean Gaussian process with the degenerate covariance function $C_{m}$ where
$\tau_{m}(\bm{s})$ $=E[\delta_{m}(\bm{s})|\Phi^{(m)}\bm{\delta}_{m}(Q^{(m)})]$
and
Finally, in the MRA-lp, the original covariance matrix $C_{0}(S_{0},S_{0})$ is approximated by $\sum_{m=0}^{M}C_{\tau_{m}}(S_{0},S_{0})$ where $S_{0}$ is a set of observed locations on $D_{0}$. Similar to the relationship between the linear projection and the predictive process, the MRA-lp is identical with the MRA in the case of $\Phi_{j_{1},\ldots,j_{m}}=\bm{\mbox{I}}_{\left|Q_{j_{1},\ldots,j_{m}}\right|}$.
In comparison with the MRA, the MRA-lp includes $\Phi_{j_{1},\ldots,j_{m}}$ which compresses the knots by taking the linear combination. The introduction of $\Phi_{j_{1},\ldots,j_{m}}$ has some advantages. First, we can separately select the number $r_{j_{1},\ldots,j_{m}}$ of knots and the rank $r^{\prime}_{j_{1},\ldots,j_{m}}$ of $\Phi_{j_{1},\ldots,j_{m}}$. Therefore, we can employ more knots than the MRA by setting the low rank of $\Phi_{j_{1},\ldots,j_{m}}$, which can lead to more efficient computation. Second, it may alleviate the selection problem of the knots. Finally, the remarkable property of $\Phi_{j_{1},\ldots,j_{m}}$ is to yield the numerical stability in calculating the inverse matrix. The MRF-lp also has these merits obtained by introducing $\Phi_{j_{1},\ldots,j_{m}}$ in contrast to the MRF.
@@ 6.1 Simulation study
Through simulation studies, we evaluated the performances of Algorithms 35. We used Algorithms 35 with $\Phi_{j_{1},\ldots,j_{m}}=\bm{\mathrm{I}}_{r_{j_{1},\ldots,j_{m}}}$ as the MRF, meaning $r_{j_{1},\ldots,j_{m}}=r_{j_{1},\ldots,j_{m}}^{\prime}$ in the MRF. The original methods in the non-Gaussian case correspond to Algorithms 4 and 5 where Algorithm 2 is not applied. Additionally, we conducted a comparison with the EnKF.
@@
Sixth, we checked the numerical stability of the MRF and MRF-lp and specified a particular scenario where the MRF-lp outperforms the MRF in terms of the MSPE. In this simulation study, we focused on the linear and Gaussian case with the Matérn covariance function with the smoothness parameter 3.5 and the Gaussian (squared exponential) covariance function. The range parameters were 1, 3, 5, and 7, and the count of iterations for calculating the ratios of the totally averaged MSPE at all of the spatiotemporal grid points was 30. When a pair of observations has a negligible ($<$ 0.05) correlation, the distance between the observations is called the effective range. For the range parameters 1, 3, 5, and 7, their effective ranges were 6.88, 2.30, 1.38, and 0.98, respectively, in the Matérn covariance function with the smoothness parameter 3.5. Similarly, in the Gaussian covariance function, they were 1.74, 1.0, 0.78, and 0.66 for the range parameters 1, 3, 5, and 7, respectively. For the MRF, we adopted $(r_{0},r_{j_{1}},r_{j_{1},j_{2}},r_{j_{1},j_{2},j_{3}},r_{j_{1},j_{2},j_{3},j_{4}})=(r_{0}^{\prime},r_{j_{1}}^{\prime},r_{j_{1},j_{2}}^{\prime},r_{j_{1},j_{2},j_{3}}^{\prime},r_{j_{1},j_{2},j_{3},j_{4}}^{\prime})=(10,10,10,2,2)$. For the MRF-lp, we set the same knot case and the same rank case, that is, $(r_{0},r_{j_{1}},r_{j_{1},j_{2}},r_{j_{1},j_{2},j_{3}},r_{j_{1},j_{2},j_{3},j_{4}})=(10,10,10,2,2)$ and $(r_{0}^{\prime},r_{j_{1}}^{\prime},r_{j_{1},j_{2}}^{\prime},r_{j_{1},j_{2},j_{3}}^{\prime},r_{j_{1},j_{2},j_{3},j_{4}}^{\prime})=(7,7,7,2,2)$ (Case 1) and $(r_{0},r_{j_{1}},r_{j_{1},j_{2}},r_{j_{1},j_{2},j_{3}},r_{j_{1},j_{2},j_{3},j_{4}})=(20,20,20,4,4)$ and $(r_{0}^{\prime},r_{j_{1}}^{\prime},r_{j_{1},j_{2}}^{\prime},r_{j_{1},j_{2},j_{3}}^{\prime},r_{j_{1},j_{2},j_{3},j_{4}}^{\prime})=(10,10,10,2,2)$ (Case 2). The other conditions were the same as the baseline of the linear and Gaussian case in Section 6.1.
Table 2: Comparative performance of the MRF and MRF-lp with respect to the MSPE under the strong spatial correlation.
Covariance Range MRF MRF-lp (Case 1) MRF-lp (Case 2)
Matérn (3.5) 1 - 1.034 1.019
3 1.041 1.084 1.022
5 1.095 1.139 1.094
7 1.121 1.199 1.116
Gaussian 1 - 1.042 1.028
3 1.107 1.138 1.067
5 1.125 1.204 1.110
7 1.172 1.246 1.138
Table 2 shows the MSPE in each case. In the same covariance case of Table 2, the larger the effective range was, the smaller the MSPE was. The large effective range means that the large-scale spatial variation is dominant. Therefore, this may be because the MRA and the MRA-lp, respectively, extend the predictive process (Banerjee et al.,, 2008) and the linear projection (Banerjee et al.,, 2013) which are originally effective for fitting the large-scale spatial variation. Moreover, we cannot conduct the MRF in the cases of the range parameter 1 because the output of the MRA was unstable and included outliers due to the numerical instability of the calculation related to the inverse matrix of $\widehat{V}_{j_{1},\ldots,j_{m}}^{m}$ ($m=0,\ldots,M$) in the MRA implemented in the MRF, that is, Algorithm 2 in Algorithm 3 with $\Phi_{j_{1},\ldots,j_{m}}=\bm{\mathrm{I}}_{r_{j_{1},\ldots,j_{m}}}$. This numerical stability is evaluated by the condition number which means the ratio of the largest $\sigma_{l}$ and the smallest $\sigma_{s}$ eigenvalues of the positive definite matrix (see Dixon,, 1983) and evaluates how well the positive definite matrix is conditioned. If a positive definite matrix is ill-conditioned, the calculation of the inverse matrix may be unstable with the propagation of round-off errors due to the finite precision arithmetic. The condition number closer to 1 indicates better numerical stability. We compared the condition number of $\widehat{V}_{j_{1},\ldots,j_{m}}^{m}$ ($m=0,\ldots,M$) between the MRF and the MRF-lp. This simulation study is similar to those in Section 3.2 of Banerjee et al., (2013) and Section 4.3 of Hirano, (2021). We confirmed that as the number $r_{j_{1},\ldots,j_{m}}$ of knots, $m$, the rank $r_{j_{1},\ldots,j_{m}}^{\prime}$ of $\Phi_{j_{1},\ldots,j_{m}}$, the smoothness parameter, and the effective range increased, the condition numbers of $\widehat{V}_{j_{1},\ldots,j_{m}}^{m}$ tended to increase. However, our proposed MRF-lp empirically showed smaller conditional numbers than those of the MRF. This may be because the MRA-lp introduces $\Phi_{j_{1},\ldots,j_{m}}$, the minor eigenvalues of $V_{j_{1},\ldots,j_{m}}^{m}$ can be removed in the selection of $\Phi_{j_{1},\ldots,j_{m}}$ (see Section 3.3 for details), and Proposition 1 holds. These results are not reported here. In the cases of the range parameter 1, their effective ranges were large relative to the size of the domain $[0,1]^{2}$. However, in the case of the exponential covariance function, we can conduct the MRF for the appropriate knots even when the effective range is greater than 100. This suggests that both the large smoothness parameter and the large effective range degrade the numerical stability of the MRF. The MRF-lp can select more knots and improve the MSPE under the strong spatial correlation with the large smoothness parameter and effective range. On the other hand, unlike the MRF-lp, it is difficult for the MRF to improve the MSPE to some extent by adding more knots because increasing knots often causes an infeasible result due to the high condition number of $\widehat{V}_{j_{1},\ldots,j_{m}}^{m}$. Actually, in the case of the range parameter 1, we needed to reduce the number of knots of the MRF in order to implement the MRF normally. In this case, most of the feasible trials of the MRF produced larger MSPEs than those of the MRF-lp (Case 1), and all of their MSPEs by the MRF were larger than a feasible MSPE by the MRF-lp. It is suggested that it is difficult to set the appropriate knot size in the MRF under the strong spatial correlation. We also obtained similar results for the nonlinear and non-Gaussian case.
@@ 7 Conclusion and future studies
This study proposed the MRF-lp, which can quickly compute the filtering mean for linear Gaussian state-space models. Furthermore, we have extended our proposed method to nonlinear and non-Gaussian state-space models on the basis of Zilber and Katzfuss, (2021) and Jurek and Katzfuss, (2022). The block sparse structure resulting from the MRA-lp makes the MRF-lp the efficient filtering algorithm. The MRF-lp can be regarded as an extension of the MRA-lp to the spatio-temporal datasets and a generalization of the MRF. By increasing $r_{j_{1},\ldots,j_{m}}$ and/or $r^{\prime}_{j_{1},\ldots,j_{m}}$ adequately, the approximation accuracy of the filtering mean can be improved while maintaining the relatively small computational time. Our proposed MRF-lp has the advantage of adjusting $r_{j_{1},\ldots,j_{m}}$ and $r^{\prime}_{j_{1},\ldots,j_{m}}$ separately, unlike the MRF where $r_{j_{1},\ldots,j_{m}}=r^{\prime}_{j_{1},\ldots,j_{m}}$. In the simulation studies and real data analysis, the MRF-lp was generally more efficient than the MRF in terms of approximation accuracy and computational time. Moreover, the MRF-lp empirically improved the numerical stability of the MRF.
There are some issues to be tackled in the future. First, Jurek and Katzfuss, (2021, 2022) proposed the fast computation algorithm of Rao–Blackwellized particle filter (Doucet and Johansen,, 2011) to estimate parameters in state-space models (e.g., range parameter in the covariance function). Similarly, our proposed MRF-lp would be extended to the particle version. Second, it is difficult for the extended Kalman filter to capture the strongly nonlinear evolution. Based on Chakraborty and Katzfuss, (2022), combining the MRF-lp with the unscented Kalman filter is interesting. Third, similar to Jurek and Katzfuss, (2023), the fast smoothing algorithm by using the MRA-lp is considered. As shown in Figure 6 of Section 6.2, the filtering distribution degenerates when few observations are obtained at early time points. However, the smoothing distribution may provide a plausible prediction surface. Fourth, the MRF-lp has a lot of tuning parameters such as the highest resolution $M$, the partition number $J_{m}$ of each subregion, the number $r_{j_{1},\ldots,j_{m}}$ of knots, and the rank $r_{j_{1},\ldots,j_{m}}^{\prime}$ of $\Phi_{j_{1},\ldots,j_{m}}$. A practical guideline for choosing them would be helpful. Finally, the general Vecchia approximation proposed by Katzfuss and Guinness, (2021) includes some existing fast computation methods used in spatial statistics and has shown high usefulness in a variety of situations (e.g., Katzfuss et al., 2020a, ; Zilber and Katzfuss,, 2021; Katzfuss et al.,, 2022; Jurek and Katzfuss,, 2022, 2023). It is left for a future study to explore the relationship between the MRF-lp and the fast filtering method in Jurek and Katzfuss, (2022), which adopts a hierarchical Vecchia approximation.